OpenStudy (anonymous):

Does anyone know how to solve U/5+U/20+U/6=1

OpenStudy (jdoe0001):

between 5 20 and 6 can you give me a number that all those 3 guys can divide it?

OpenStudy (anonymous):

I cant think of a number that can be divided into each number

OpenStudy (jdoe0001):

\(\bf \square \div 5\qquad \square \div 20 \qquad \square \div 6\)

OpenStudy (anonymous):

Wait I wrote the problem wrong .. It's suppose to be \[\frac{ U }{ 5 } + \frac{ U }{ 10} + \frac{ U }{ 6 } = 1\]

OpenStudy (anonymous):

so would it be 30

OpenStudy (jdoe0001):

\(\bf \cfrac{u}{5}+\cfrac{u}{10}+\cfrac{u}{6}=1\implies \cfrac{u}{5}+\cfrac{u}{10}+\cfrac{u}{6}=\cfrac{1}{1} \)

OpenStudy (jdoe0001):

so your denominators are 5, 10, 6 and 1 the LCD will be the a number that's divisible by all 4 of those guys, I initially didn't include "1" because any number is divisible by 1

OpenStudy (jdoe0001):

well, let's try 30 \(\bf 30 \div 5 = 6\qquad \qquad 30 \div 10 = 3\qquad \qquad 30 \div 6 = 5\)

OpenStudy (anonymous):

So saying That all of the denominators have to be the same?

OpenStudy (jdoe0001):

well, let us use the LCD, there's an easier method but .... let's see this one so \(\bf \cfrac{u}{5}+\cfrac{u}{10}+\cfrac{u}{6}=\cfrac{1}{1} \implies \cfrac{u}{5}+\cfrac{u}{10}+\cfrac{u}{6}-\cfrac{1}{1}=0\\ \cfrac{\square+\square+\square-\square}{30}\)

OpenStudy (jdoe0001):

do you know what goes atop?

OpenStudy (anonymous):

0?

OpenStudy (jdoe0001):

\(\bf \cfrac{\square+\square+\square-\square}{30}=0\) rather

OpenStudy (anonymous):

So I have to find numbers that will add and subtract together and be divided by 30 to equal 0

OpenStudy (jdoe0001):

\(\bf \cfrac{u}{5}+\cfrac{u}{10}+\cfrac{u}{6}-\cfrac{1}{1}=0 \implies \cfrac{\left(\frac{30}{5}\right)u+\left(\frac{30}{10}\right)u+\left(\frac{30}{6}\right)u-\left(\frac{30}{1}\right)u}{30}\)

OpenStudy (jdoe0001):

when adding fractions, you use LCD, divide it by the fraction's denominator's and multiply the quotient by the numerator, that goes up there, in the resulting fraction

OpenStudy (jdoe0001):

hmm, the last one has no "u"... lemme fix that

OpenStudy (jdoe0001):

\(\bf \cfrac{u}{5}+\cfrac{u}{10}+\cfrac{u}{6}-\cfrac{1}{1}=0 \implies \cfrac{\left(\frac{30}{5}\right)u+\left(\frac{30}{10}\right)u+\left(\frac{30}{6}\right)u-\left(\frac{30}{1}\right)}{30}\)

OpenStudy (jdoe0001):

also forgot to equate to 0, ohh well, anyhow the last fraction is really unnecessary, but anyhow

OpenStudy (anonymous):

So I wouldn't have to put \[\frac{ 30 }{ 1 }\]

OpenStudy (jdoe0001):

well, I could have written it as \(\bf \cfrac{u}{5}+\cfrac{u}{10}+\cfrac{u}{6}=1 \implies \cfrac{\left(\frac{30}{5}\right)u+\left(\frac{30}{10}\right)u+\left(\frac{30}{6}\right)u}{30} = 1\)

OpenStudy (jdoe0001):

and once you get your numerator above, you can always cross-multiply

OpenStudy (jdoe0001):

the result will be the same, just less clutter

OpenStudy (anonymous):

so what would be the result

OpenStudy (jdoe0001):

what did you get in the numerator?

OpenStudy (anonymous):

Will there be 3 different numbers?

OpenStudy (jdoe0001):

should, yes

OpenStudy (anonymous):

So I have 2, 3, 2

OpenStudy (jdoe0001):

30/5 = 2? 30/6 = 2? \(\bf \cfrac{\left(\color{red}{\frac{30}{5}}\right)u+\left(\color{red}{\frac{30}{10}}\right)u+\left(\color{red}{\frac{30}{6}}\right)u}{30} = 1\)

OpenStudy (anonymous):

My bad .. I meant .. 6 and 5

OpenStudy (jdoe0001):

\(\bf \cfrac{u}{5}+\cfrac{u}{10}+\cfrac{u}{6}=1 \implies \cfrac{\left(\frac{30}{5}\right)u+\left(\frac{30}{10}\right)u+\left(\frac{30}{6}\right)u}{30} = 1 \\ \cfrac{6u+3u+5u}{30}=1\implies \cfrac{14u}{30} = 1\) now solve for "u"

OpenStudy (anonymous):

I get huge decimal when I solve for u

OpenStudy (jdoe0001):

well, don't use a decimal :), leave it as fraction

OpenStudy (anonymous):

So \[\frac{ 14u }{ 30 } = 1 \] is the answer

OpenStudy (jdoe0001):

no, you'd need to solve it for "u"

OpenStudy (anonymous):

How?

OpenStudy (jdoe0001):

or isolate "u" that is

OpenStudy (jdoe0001):

http://www.youtube.com/watch?v=_A8lLbZCrlw

OpenStudy (anonymous):

so would you do the reciprocal?

OpenStudy (jdoe0001):

you 'd multiply both sides by 30, then divide both by 14

OpenStudy (anonymous):

I got 2.14

OpenStudy (jdoe0001):

if you use decimals, yes

OpenStudy (anonymous):

Okay. we do .